# Compact Visual Data Representation for Green Multimedia - A Human Visual System Perspective

### v1

@@ III-A1 Video Coding Standards
More specifically, 67 modes have been designed for accommodating diverse video content textures [28]. This granularity enhances prediction accuracy and reduces residual signaling burdens. Meanwhile, the adoption of multiple reference lines and the Cross-Component Linear Model (CCLM) predictions further leverage inter-channel redundancies for chroma intra coding [29, 30, 31], showcasing significant efficiency in handling color components. Moreover, the VVC and AVS3 standards incorporate a range of sophisticated intra prediction tools tailored for processing non-natural video content. These include intra block copy [32, 33], palette mode [34], and string copy prediction [35]. On the inter prediction front, the standards excel at reducing temporal redundancies. To handle the complex motions in real-world scenes, VVC and AVS3 have integrated affine motion compensation techniques [36, 37], designed to handle non-translational movements like zooming and rotation. The rise in video resolution further complicates MV signaling, prompting the development of sophisticated MV coding and refinement strategies. These include adaptive motion vector resolution (AMVR)[38], history-based motion vector prediction (HMVP)[39], advanced motion vector prediction (AMVP)[40], bi-directional optical flow (BDOF)[41], merge mode with motion vector difference (MMVD)[42], and decoder-side motion vector refinement (DMVR)[43]. For residual representation in both intra and inter prediction, trellis-coded quantization (TCQ)[44] is employed, where coefficients within a coding block are mapped onto a trellis graph to determine the path with the lowest rate-distortion cost, effectively minimizing redundancies. These standards support multiple discrete cosine or sine-based transform (DCT/DST) cores, which help to more efficiently concentrate residual energies in VVC[45] and AVS3 [46]. Historically, overlapped block motion compensation has been used to enhance the accuracy of inter prediction [47]. Finally, the elimination of statistical redundancy is achieved through context-adaptive binary arithmetic coding (CABAC) [48], which merges adaptive binary arithmetic coding with context modeling for statistical redundancy removal. Recently, the Enhanced Compression Model (ECM), a further enhancement of the VVC, has undergone extensive studies with a variety of advanced coding tools. These tools encompass various coding modules such as intra/inter prediction, aiming to further enhance its performance [49, 50, 51, 52, 53, 54].
@@ III-A2 End-to-end Video Coding Technologies
The end-to-end (E2E) coding framework serves as an alternative solution for compression. Unlike traditional methods that often involve separate components for different stages of compression, E2E frameworks integrate all coding processes from input to output into a unified system. This integration allows the system to learn and optimize the compression pipeline holistically, using deep learning techniques to adaptively manage data based on content complexity and desired fidelity levels. By training these systems directly on the target data, they can potentially achieve higher efficiency and better quality compared to modular approaches. In exploring end-to-end visual data compression, a notable pioneer contribution is made by Ballé et al., who designed an end-to-end architecture that includes a decorrelation layer known as Generalized Divisive Normalization (GDN), and replaced traditional step quantization with uniform noise addition during training to highlight the advantages of end-to-end methods [68, 69]. Subsequent studies have typically utilized the upper bound entropy of latent codes as a surrogate for the rate term in rate-distortion optimization. Ballé et al. enhanced their architecture with a hyper-prior network that aids in decoding by leveraging side information extracted from latent codes, presuming a Gaussian distribution [70]. This approach was further refined by Cheng et al., who employed a Gaussian mixture model instead of a simple Gaussian for a more precise entropy estimation [71]. Chang et al. proposed a framework that utilizes deep structural and textural information to enhance compression versatility [72]. Recently, a hybrid video compression framework based on the ECM [73] has been proposed to integrate deep learning techniques, achieving significant BD-rate savings for Y, U, and V components under random access configuration. The field of end-to-end video compression has also seen significant advancements. Lu et al. developed deep video compression (DVC), which is the first comprehensive model for video compression that optimizes nearly all aspects of the framework, including the use of a hyper-prior model for intra-frame compression and an optical flow network for effective motion estimation and compensation [74]. They later enhanced this model with an adaptive quantization layer to support variable bitrate coding [75]. Building on traditional video coding techniques, Yang et al. introduced a strategy that incorporates hierarchical quality sub-models, yielding results competitive with the fastest modes of the HEVC standard [76]. Recently, Li et al. introduced a deep contextual video compression method named DCVC [77], which shifts from the traditional residual coding strategy to a conditional coding framework. Instead of using subtraction to eliminate temporal redundancy, DCVC treats temporal prediction as conditions, enabling the codec to mine the potential dependencies implicitly. Subsequently, Sheng et al. introduced DCVC-TCM [78], which enhances DCVC by integrating multi-scale temporal predictions throughout the encoding process and using them for entropy modeling. Li et al. expanded on this with DCVC-HEM [79], adding a latent prior to further optimize temporal correlation. By increasing the diversity of contexts in DCVC, DCVC-DC was proposed to achieve superior coding performance compared with VVC [80]. Recently, DCVC-FM [81] adopted feature modulation to support a wider quality range and maintain performance under long prediction chains, achieving significant bitrate savings and computational efficiency improvements. NVC-1B [82] is the first neural video coding model with over 1 billion parameters, demonstrating significant improvements in video compression performance and setting a new state-of-the-art in compression efficiency.
@@ III-A3 Perceptual Compression
In the context of green multimedia, perceptual coding offers an energy-efficient approach by leveraging the HVS to focus computational resources on areas that significantly impact perceived quality. By reducing the processing complexity in regions less critical to visual experience, perceptual coding techniques can decrease the amount of data required for transmission and storage, thus reducing bandwidth and energy consumption across the multimedia pipeline without compromising user experience. Perceptual compression is a long-standing research topic that has attracted numerous attentions [83, 84]. The fundamental principle is to optimize compression efficiency by simulating visual perception and exploiting visual redundancies. While PSNR remains a widely used metric due to its simplicity and friendly hardware implementations, various visual sensitivity models have been developed to better align with perceptual quality requirements. Hence, many visual sensitivity models have been proposed by studying the characteristics of visual distortions. Specifically, the feature-similarity (FSIM) index has been introduced, which considers the contribution of low-level features perceived by HVS [85]. The nature of compression artifacts that affect HVS also necessitates research to characterize and mitigate their impact on perceptual quality [86]. The Video Multimethod Assessment Fusion (VMAF) metric [87] offers a highly efficient full-reference video quality assessment method suitable for a wide range of video coding scenarios. Recent advancements reveal that deep features from neural networks can significantly enhance perceptual similarity assessments, proposing a paradigm shift in how perceptual quality is evaluated in compression [88]. The utilization of visual Just Noticeable Difference (JND) modeling is also one of the research directions of perceptual compression [89, 90, 91, 92]. These approaches leverage the fact that human eyes cannot detect changes below the JND threshold around a pixel, due to their inherent spatial and temporal sensitivity as well as the masking properties [93]. Visual Attention (VA), which is also referred to as saliency, is a cognitive process that directs the eye’s focus toward Regions of Interest (ROI) to capture finer details. This mechanism can also be leveraged to effectively advance the perceptual compression [94, 95]. Furthermore, Chen et al. introduced a novel approach with the proxy network to optimize the end-to-end compression model against quantitative perceptual performance [96]. Rouis et al. adapts the Lagrangian multiplier based on perceptual criteria within the HEVC standard, aiming to enhance coding efficiency by focusing on visually significant areas of the video [97]. The perceptually tuned ECM has also been studied [98], which aims to maximise the perceived quality. Additionally, spatial domain pre-processing is a popular strategy where blur intensity is tailored to make less significant regions more easily encodable. To make preprocessing compatible with diverse codecs, Chadha et al. proposed a rate-aware perceptual-oriented framework comprised of dedicated perceptual loss and rate loss to achieve single-pass operation before video coding [99]. Arai et al. [100] proposes a novel GOP-based method that enhances rate-distortion performance in terms of PSNR and other quality metrics by jointly optimizing a preprocessing model with the video compression model. One of the primary benefits of these preprocessing techniques is their compatibility with any standard codec, eliminating the need for modifications. Nonetheless, since the quality control and encoding processes are independent, the gains from preprocessing are generally considered a lower bound on the potential improvements in perceptual compression. Moreover, combining visual encoding with existing standards allows for seamless integration into current systems without additional hardware or decoder updates. In general, video coding standards, including H.264/AVC, AVS2, HEVC, AVS3, and VVC, support flexible encoding parameters across different regions, enabling the assignment of varied quality parameters in a frame. This approach allows for finer quality distinctions, with critical visual areas encoded at higher quality due to lower QPs.
@@ III-C Unified Representation for Dynamic Tasks
Unified representation frameworks aim to generate a single compressed representation suitable for multiple dynamic tasks. By enabling adaptable data streams that adjust to task-specific requirements, these frameworks eliminate redundant data processing and transmission. In green multimedia, this approach enhances energy efficiency by dynamically aligning resource usage with task needs, reducing the overall data load and processing requirements. This is aligned with the mechanism of the human brain, where the input signals across different regions and hierarchical levels can be adaptively integrated according to the demands of various tasks. However, the traditional coding targets at the reconstruction of only feature or texture, which cannot support multiple tasks simultaneously. Regarding the diverse goals of downstream tasks, recent efforts have been devoted to scalable layered coding, including features and textures that uniquely contribute to the efficient representation of visual information from different aspects. For example, in [154], the base layer accounts for the compact feature and the enhancement layer reconstructs the texture. Lin et al. proposed to divide the bitstream into semantic, structure, and texture layers to achieve both machine and human vision [155]. These integrated approaches, though still far from the true mechanism of HVS which has not been well understood, still offers several unique benefits. For example, the compact representation of high-quality features extracted directly from the original texture could largely enhance the analytical capabilities. Second, on the decoding end, features can be processed directly, significantly reducing the computational load compared to standard video decoding and separate feature extraction. Thirdly, the synergy between feature and texture coding improves the overall efficiency of the joint bitstream, since the redundancy among the two representations has been already removed.
@@
Dynamic green operations not only promote significant energy savings but also enhance the adaptability and efficiency of multimedia systems. By tailoring computational efforts to task needs, these operations optimize resource allocation, minimizing unnecessary power consumption and extending the operational lifespan of devices. Furthermore, the dynamic adaptation of multimedia processing mirrors the flexibility and efficiency of biological systems, such as the human brain, which optimizes its functions based on environmental demands and internal states [161]. By integrating such biologically-inspired approaches, dynamic green operations pave the way for more sustainable and intelligent multimedia technologies. Specifically, to further elaborate on the potential of scalable coding schemes, Choi et al. presented an end-to-end learned image codec optimized for both human viewing and machine vision tasks, which utilizes a layered approach where the base layer is dedicated to simple tasks and additional layers enhance capability for more complex tasks [162]. To jointly leverage the priors from the external database and face sketch, an ultra-low bitrate face compression was developed to reconstruct perceptually meaningful results while providing analytical performance [103]. Recently, Shen et al. proposed a versatile neural video coding (VNVC) framework, which not only designs highly compact representations for intermediate features but also enables their generalization in diverse downstream tasks, including reconstruction, enhancement, and analysis [163]. Ge et al. designed an innovative encoder controller for DVC that employs mode prediction and Group of Pictures (GoP) selection to adapt the pre-trained DVC decoder for various tasks [164]. Li et al. introduced a human-machine collaborative image coding framework utilizing Implicit Neural Representations (INR) [165]. This framework balances high-efficiency visual data compression for human vision with reduced information transmission for machine vision tasks. Zhang et al. proposed a Unified and Scalable Deep Image Compression (USDIC) framework that adopts a novel approach by introducing an Information Splitting Mechanism (ISM) that effectively segregates semantic and visual features for optimized machine analysis and human viewing, respectively [166].

### v2

@@ III-A1 Video Coding Standards
More specifically, 67 modes have been designed for accommodating diverse video content textures [28]. This granularity enhances prediction accuracy and reduces residual signaling burdens. Meanwhile, the adoption of multiple reference lines and the Cross-Component Linear Model (CCLM) predictions further leverage inter-channel redundancies for chroma intra coding [29, 30, 31], showcasing significant efficiency in handling color components. Moreover, the VVC and AVS3 standards incorporate a range of sophisticated intra prediction tools tailored for processing non-natural video content. These include intra block copy [32, 33], palette mode [34], and string copy prediction [35]. On the inter prediction front, the standards excel at reducing temporal redundancies. To handle the complex motions in real-world scenes, VVC and AVS3 have integrated affine motion compensation techniques [36, 37], designed to handle non-translational movements like zooming and rotation. The rise in video resolution further complicates MV signaling, prompting the development of sophisticated MV coding and refinement strategies. These include adaptive motion vector resolution (AMVR)[38], history-based motion vector prediction (HMVP)[39], advanced motion vector prediction (AMVP)[40], bi-directional optical flow (BDOF)[41], merge mode with motion vector difference (MMVD)[42], and decoder-side motion vector refinement (DMVR)[43]. For residual representation in both intra and inter prediction, trellis-coded quantization (TCQ)[44] is employed, where coefficients within a coding block are mapped onto a trellis graph to determine the path with the lowest rate-distortion cost, effectively minimizing redundancies. These standards support multiple discrete cosine or sine-based transform (DCT/DST) cores, which help to more efficiently concentrate residual energies in VVC[45] and AVS3 [46]. Historically, overlapped block motion compensation has been used to enhance the accuracy of inter prediction [47]. Finally, the elimination of statistical redundancy is achieved through context-adaptive binary arithmetic coding (CABAC) [48], which merges adaptive binary arithmetic coding with context modeling for statistical redundancy removal. Regarding encoding optimization, a multi-grained parallel solution (MPS) was proposed to optimize HEVC encoding on heterogeneous systems by leveraging massively parallel motion estimation on GPUs and adaptive parallel mode decisions on CPUs, demonstrating significant improvements in concurrent execution efficiency [49]. Recently, the Enhanced Compression Model (ECM), a further enhancement of the VVC, has undergone extensive studies with a variety of advanced coding tools. These tools encompass various coding modules such as intra/inter prediction, aiming to further enhance its performance [50, 51, 52, 53, 54, 55].
@@ III-A2 End-to-end Video Coding Technologies
The end-to-end (E2E) coding framework serves as an alternative solution for compression. Unlike traditional methods that often involve separate components for different stages of compression, E2E frameworks integrate all coding processes from input to output into a unified system. This integration allows the system to learn and optimize the compression pipeline holistically, using deep learning techniques to adaptively manage data based on content complexity and desired fidelity levels. By training these systems directly on the target data, they can potentially achieve higher efficiency and better quality compared to modular approaches. In exploring end-to-end visual data compression, a notable pioneer contribution is made by Ballé et al., who designed an end-to-end architecture that includes a decorrelation layer known as Generalized Divisive Normalization (GDN), and replaced traditional step quantization with uniform noise addition during training to highlight the advantages of end-to-end methods [69, 70]. Subsequent studies have typically utilized the upper bound entropy of latent codes as a surrogate for the rate term in rate-distortion optimization. Ballé et al. enhanced their architecture with a hyper-prior network that aids in decoding by leveraging side information extracted from latent codes, presuming a Gaussian distribution [71]. This approach was further refined by Cheng et al., who employed a Gaussian mixture model instead of a simple Gaussian for a more precise entropy estimation [72]. Chang et al. proposed a framework that utilizes deep structural and textural information to enhance compression versatility [73]. Tang et al. proposed an end-to-end image compression method that integrates graph attention and asymmetric convolutional neural networks (ACNN) to enhance local key feature effects, reduce training costs, and achieve state-of-the-art rate-distortion performance [74]. Recently, a hybrid video compression framework based on the ECM [75] has been proposed to integrate deep learning techniques, achieving significant BD-rate savings for Y, U, and V components under random access configuration. The field of end-to-end video compression has also seen significant advancements. Lu et al. developed deep video compression (DVC), which is the first comprehensive model for video compression that optimizes nearly all aspects of the framework, including the use of a hyper-prior model for intra-frame compression and an optical flow network for effective motion estimation and compensation [76]. They later enhanced this model with an adaptive quantization layer to support variable bitrate coding [77]. Building on traditional video coding techniques, Yang et al. introduced a strategy that incorporates hierarchical quality sub-models, yielding results competitive with the fastest modes of the HEVC standard [78]. Recently, Li et al. introduced a deep contextual video compression method named DCVC [79], which shifts from the traditional residual coding strategy to a conditional coding framework. Instead of using subtraction to eliminate temporal redundancy, DCVC treats temporal prediction as conditions, enabling the codec to mine the potential dependencies implicitly. Subsequently, Sheng et al. introduced DCVC-TCM [80], which enhances DCVC by integrating multi-scale temporal predictions throughout the encoding process and using them for entropy modeling. Li et al. expanded on this with DCVC-HEM [81], adding a latent prior to further optimize temporal correlation. By increasing the diversity of contexts in DCVC, DCVC-DC was proposed to achieve superior coding performance compared with VVC [82]. Recently, DCVC-FM [83] adopted feature modulation to support a wider quality range and maintain performance under long prediction chains, achieving significant bitrate savings and computational efficiency improvements. NVC-1B [84] is the first neural video coding model with over 1 billion parameters, demonstrating significant improvements in video compression performance and setting a new state-of-the-art in compression efficiency. While neural network-based compression techniques can achieve high performance in terms of compression efficiency and subjective quality, their computational overhead and energy consumption still raise important sustainability concerns. Reducing this complexity without sacrificing compression effectiveness represents a critical challenge in green multimedia research. Current strategies address this trade-off by exploring model compactness techniques, such as pruning and quantization [85], which can dramatically lower computational requirements, and by leveraging knowledge distillation to transfer learned features into more lightweight models [86]. Beyond these approaches, innovative transforms and hybrid architectures that fuse traditional coding techniques with learned components offer the potential to preserve much of the efficiency gain from neural models while limiting their energy footprint [87]. As the field continues to mature, the development of even more sophisticated complexity-control mechanisms—supported by new theoretical insights and hardware acceleration strategies—will be essential. By prioritizing methods that harmonize compression efficiency with low energy consumption, future research can help ensure that advanced compression technologies remain both environmentally responsible and technically competitive.
@@ III-A3 Perceptual Compression
In the context of green multimedia, perceptual coding offers an energy-efficient approach by leveraging the HVS to focus computational resources on areas that significantly impact perceived quality. By reducing the processing complexity in regions less critical to visual experience, perceptual coding techniques can decrease the amount of data required for transmission and storage, thus reducing bandwidth and energy consumption across the multimedia pipeline without compromising user experience [88, 89, 90]. Perceptual compression is a long-standing research topic that has attracted numerous attentions [91, 92]. The fundamental principle is to optimize compression efficiency by simulating visual perception and exploiting visual redundancies. While PSNR remains a widely used metric due to its simplicity and friendly hardware implementations, various visual sensitivity models have been developed to better align with perceptual quality requirements. Hence, many visual sensitivity models have been proposed by studying the characteristics of visual distortions. Specifically, the feature-similarity (FSIM) index has been introduced, which considers the contribution of low-level features perceived by HVS [93]. The nature of compression artifacts that affect HVS also necessitates research to characterize and mitigate their impact on perceptual quality [94]. The Video Multimethod Assessment Fusion (VMAF) metric [95] offers a highly efficient full-reference video quality assessment method suitable for a wide range of video coding scenarios. Specifically, Luo et al. have designed a VMAF oriented perceptual coding Based on piecewise metric coupling, which achieves promising bit savings while better aligned with HVS [96]. Recent advancements reveal that deep features from neural networks can significantly enhance perceptual similarity assessments, proposing a paradigm shift in how perceptual quality is evaluated in compression [97]. The utilization of visual Just Noticeable Difference (JND) modeling is also one of the research directions of perceptual compression [98, 99, 100, 101]. These approaches leverage the fact that human eyes cannot detect changes below the JND threshold around a pixel, due to their inherent spatial and temporal sensitivity as well as the masking properties [102]. Visual Attention (VA), which is also referred to as saliency, is a cognitive process that directs the eye’s focus toward Regions of Interest (ROI) to capture finer details. This mechanism can also be leveraged to effectively advance the perceptual compression [103, 104, 90]. In particular, Mazumdar et al. proposed a saliency-based perceptual compression technique named Vignette, which can cooperate with off-the-shelf compression codecs to improve cloud storage capacity and reduce power consumption in mobile devices [90]. Furthermore, Chen et al. introduced a novel approach with the proxy network to optimize the end-to-end compression model against quantitative perceptual performance [105]. Rouis et al. adapts the Lagrangian multiplier based on perceptual criteria within the HEVC standard, aiming to enhance coding efficiency by focusing on visually significant areas of the video [106]. For perceptually compressed VR videos, the video decoder architecture named mVDO was proposed, which maps peripheral video regions to decoder cores in multiple frequency domains to achieve better energy efficiency [107]. The perceptually tuned ECM has also been studied [108], which aims to maximise the perceived quality. Additionally, spatial domain pre-processing is a popular strategy where blur intensity is tailored to make less significant regions more easily encodable. To make preprocessing compatible with diverse codecs, Chadha et al. proposed a rate-aware perceptual-oriented framework comprised of dedicated perceptual loss and rate loss to achieve single-pass operation before video coding [109]. Arai et al. [110] proposes a novel GOP-based method that enhances rate-distortion performance in terms of PSNR and other quality metrics by jointly optimizing a preprocessing model with the video compression model. One of the primary benefits of these preprocessing techniques is their compatibility with any standard codec, eliminating the need for modifications. Nonetheless, since the quality control and encoding processes are independent, the gains from preprocessing are generally considered a lower bound on the potential improvements in perceptual compression. Moreover, combining visual encoding with existing standards allows for seamless integration into current systems without additional hardware or decoder updates. In general, video coding standards, including H.264/AVC, AVS2, HEVC, AVS3, and VVC, support flexible encoding parameters across different regions, enabling the assignment of varied quality parameters in a frame. This approach allows for finer quality distinctions, with critical visual areas encoded at higher quality due to lower QPs. Furthermore, exploring non-PSNR distortion measures, such as SSIM or VMAF, in conjunction with ROI techniques presents a promising avenue for future research, particularly in the context of power saving. While these methods are known to enhance visual quality, their potential for power reduction remains underexplored and warrants further investigation. Future researchs can explore the relationship between perceptual video coding and power efficiency, examining how these advanced distortion measures can contribute to energy savings at various video processing stages.
@@ III-C Unified Representation for Dynamic Tasks
Unified representation frameworks aim to generate a single compressed representation suitable for multiple dynamic tasks. By enabling adaptable data streams that adjust to task-specific requirements, these frameworks eliminate redundant data processing and transmission. In green multimedia, this approach enhances energy efficiency by dynamically aligning resource usage with task needs, reducing the overall data load and processing requirements. This is aligned with the mechanism of the human brain, where the input signals across different regions and hierarchical levels can be adaptively integrated according to the demands of various tasks. However, the traditional coding targets at the reconstruction of only feature or texture, which cannot support multiple tasks simultaneously. Regarding the diverse goals of downstream tasks, recent efforts have been devoted to scalable layered coding, including features and textures that uniquely contribute to the efficient representation of visual information from different aspects. For example, in [164], the base layer accounts for the compact feature and the enhancement layer reconstructs the texture. Lin et al. proposed to divide the bitstream into semantic, structure, and texture layers to achieve both machine and human vision [165]. Although these integrated approaches are still far from the actual mechanism of HVS, which has not been well understood, they still offer several unique benefits. For example, the compact representation of high-quality features extracted directly from the original texture could largely enhance the analytical capabilities. Furthermore, at the decoding end, features can be processed directly from the bitstream syntax elements [166], such as motion vectors or coding modes, which are part of the video coding standard. This strategy significantly reduces the computational load compared to standard video decoding and separate feature extraction, potentially enhancing the performance of downstream tasks such as object detection. Thirdly, the synergy between feature and texture coding improves the overall efficiency of the joint bitstream, since the redundancy among the two representations has been already removed.
@@
Dynamic green operations not only promote significant energy savings but also enhance the adaptability and efficiency of multimedia systems. By tailoring computational efforts to task needs, these operations optimize resource allocation, minimizing unnecessary power consumption and extending the operational lifespan of devices. Furthermore, the dynamic adaptation of multimedia processing mirrors the flexibility and efficiency of biological systems, such as the human brain, which optimizes its functions based on environmental demands and internal states [172]. By integrating such biologically-inspired approaches, dynamic green operations pave the way for more sustainable and intelligent multimedia technologies. Specifically, to further elaborate on the potential of scalable coding schemes, Choi et al. presented an end-to-end learned image codec optimized for both human viewing and machine vision tasks, which utilizes a layered approach where the base layer is dedicated to simple tasks and additional layers enhance capability for more complex tasks [173]. To jointly leverage the priors from the external database and face sketch, an ultra-low bitrate face compression was developed to reconstruct perceptually meaningful results while providing analytical performance [113]. Notably, Yi et al. proposes a task-driven video compression framework optimized for both human and machine vision by incorporating multi-scale motion estimation, multi-frame feature fusion, and reference-based in-loop filters, achieving a balance between signal and semantic fidelity as well as superior rate-distortion and rate-accuracy performance [174]. Recently, Shen et al. proposed a versatile neural video coding (VNVC) framework, which not only designs highly compact representations for intermediate features but also enables their generalization in diverse downstream tasks, including reconstruction, enhancement, and analysis [175]. Ge et al. designed an innovative encoder controller for DVC that employs mode prediction and Group of Pictures (GoP) selection to adapt the pre-trained DVC decoder for various tasks [176]. Li et al. introduced a human-machine collaborative image coding framework utilizing Implicit Neural Representations (INR) [177]. This framework balances high-efficiency visual data compression for human vision with reduced information transmission for machine vision tasks. Zhang et al. proposed a Unified and Scalable Deep Image Compression (USDIC) framework that adopts a novel approach by introducing an Information Splitting Mechanism (ISM) that effectively segregates semantic and visual features for optimized machine analysis and human viewing, respectively [178].